What is the break-even tax rate between a taxable corporate bond yielding 7 percent and a comparable nontaxable municipal bond yielding 5 percent is?

*Please show all work..*
Answer #1

corporate bond yield*(1-tax rate)=municipal bond yield

7*(1-tax rate)=5

1-tax rate=(5/7)

tax rate=1-(5/7)

=28.57%(Approx).